# Motivic Steenrod operations at the characteristic via infinite ramification

### v1

@@ 1. Operations? Again?
At this point, the algebraic geometer might be rightly annoyed — it seems that all the topologist does is to spoil algebro-geometric dreams using these hieroglyphic-like symbols. Then comes Voevodsky. The motivic Steenrod operations on $-\ell$ motivic cohomology in characteristic zero [Voe03b] lie at the heart of the proof of the Bloch-Kato conjectures [Voe03a, Voe11]. In fact, the shape of these operations is at the heart of many issues of a good theory of motives and motivic homotopy. Another notable result is Feng’s thesis [Fen20] where he used Steenrod operations on étale cohomology to settle a conjecture of Tate’s on the Artin-Tate pairing on Brauer groups away from the residue characteristic.
@@ 1.1. Operations and endomorphisms
Let us now explain how motivic cohomology operations work. Let $p$ be a prime, $F$ a field, and let $H\mathbb{F}_{p,F}$ be the motivic spectrum representing mod-$p$ motivic cohomology in Morel–Voevodsky’s category $\mathrm{SH}_{F}$ of motivic spectra over $F$; this motivic spectrum could be built as Bloch’s cycle complex [Blo86] or via Voevodsky’s theory of finite correspondences [VSF00]. As explained in the introduction to [HKO17], there are three related, but a priori different algebra of bigraded operations in motivic homotopy theory summarized in the following zig-zag:
@@
    $\mathcal{M}_{F}^{\star,\star}$ is the algebra of bistable operations in motivic cohomology: natural transformations $H_{\mathrm{mot}}^{p}(-;\mathbb{Z}/\ell(q))\rightarrow H_{\mathrm{mot}}^{p+r}(-;\mathbb{Z}/\ell(q+s))$ defined on $\mathrm{Sm}_{F}$ commuting with both cohomological shifts and Tate twists.
@@
    Finally, $\mathcal{A}_{F}^{\star,\star}$ is thus far only defined if $p$ is invertible in $F$. It is the subalgebra of $\mathcal{M}_{F}^{\star,\star}$ generated by power operations as constructed by Voevodsky in [Voe03b] [^1] Of note are also Brosnan’s Steenrod operations for Chow groups [Bro03] built in a similar manner to Voevodsky. ; nowadays we know that these operations arise from the norm structure on motivic cohomology in the sense of [BH21a]; see [BH21a, Example 7.25] for details.
@@
Ultimately, the construction of power operations rely on the classifying stack $B\Sigma_{n}$ where $\Sigma_{n}$ is the symmetric group on $n$-letters. In effect, Voevodsky in [Voe03b] constructs the $n$-th total power operation as the dotted arrow in the following diagram of spectra [^2] We employ the following notation: for any scheme $S$ we have an adjunction $\sigma^{\infty}:\mathrm{Sh}_{\mathrm{Nis},\mathbb{A}^{1}}(\mathrm{Sm}_{S};\mathrm{Sp})\rightleftarrows\mathrm{SH}_{S}:\omega^{\infty}$. The domain of $\sigma^{\infty}$ is the $\infty$-category of $\mathbb{A}^{1}$-invariant Nisnevich sheaves, also commonly referred to as “$S^{1}$-spectra” in the literature. We will also use the further adjunction $\Sigma^{\infty}_{\mathbb{T}}:\mathrm{Sh}_{\mathrm{Nis},\mathbb{A}^{1}}(\mathrm{Sm}_{S};\mathrm{Ani})_{\star}\rightleftarrows\mathrm{SH}_{S}:\Omega^{\infty}_{\mathbb{T}},$ between pointed $\mathbb{A}^{1}$-invariant Nisnevich sheaves of anima and $\mathrm{SH}_{S}$.
@@
The problem is that $B\Sigma_{p}$ is $\mathbb{A}^{1}$-contractible in characteristic $p>0$, hence one cannot proceed in this manner. In principle, one needs to replace $B\Sigma_{p}$ by the stack $B\mu_{p}$ and one wants a “total $\mu_{p}$-power operation”: $\omega^{\infty}H\mathbb{F}_{p}(X)\rightarrow\omega^{\infty}H\mathbb{F}_{p}(X\times B\mu_{p})$ out of which one can extract elements of $\mathcal{A}^{\star,\star}$. One approach is to endow a “flat norm” structure on $H\mathbb{F}_{p}$, extending the étale norms of [BH21a]. We do not yet have access to this technology.
@@
In particular, $H_{\mathrm{mot}}^{0}(\mathbb{F}_{p},\mathbb{F}_{p}(\star))=0$ unless =0$\star$=0⋆ = 0. Part of the content of Theorem 1.1 is that even though $\widetilde{\tau}_{p}$-multiples of the Bockstein and power operations are killed by (1), those without such multiples survive and retains their independence relations.
@@
1.2. Relations to [Pri20] and [FS]
The goal of this paper is to construct power operations on mod-$p$ motivic cohomology of $\mathbb{F}_{p}$-schemes, satisfying the analogs of the relations in topology and $-\ell$ motivic cohomology. Our approach is to directly construct a graded-$\mathbb{E}_{1}$-map
@@ 1.2. Relations to [Pri20] and [FS]
The work of Primozic [Pri20], relying on the work of Frankland–Spitzweck [FS] constructs power operations $\mathrm{P}^{i}:H_{\mathrm{mot}}^{m}(-;\mathbb{F}_{p}(n))\rightarrow H_{\mathrm{mot}}^{m+2i(p-1)}(-;\mathbb{F}_{p}(n+i(p-1))$ which satisfy the Adem and Cartan formulas on the Chow line diagonal. Let us explain how this is done.
@@
Over fields where $p$ is invertible, this map is an equivalence by [Voe03b, HKO17], but this remains unknown over characteristic $p>0$ fields. Since we do not know how exactly the dual Steenrod algebra looks like, we cannot really define the power operations by (re)taking duals. Nonetheless, the main result of [FS] shows that $\Psi$ is a split monomorphism over any field. This suggests that there is some hope in producing power operations by taking duals.
@@ 1.3.1. Syntomic realizations
where $H\mathbb{F}_{p}^{\mathrm{syn}}$ denotes the motivic spectrum that represents the mod-$p$ syntomic cohomology (see Proposition 4.13). From this, we obtain mod-$p$ power operations $\mathrm{P}^{i}$ acting on the mod-$p$ syntomic cohomology, and in particular an algebra $\mathcal{A}^{\star,\star}_{\mathrm{syn}}$ acting on $H\mathbb{F}_{p}^{\mathrm{syn}}$ that is generated by the $\mathrm{P}^{i}$ and the Bockstein $\beta$ modulo the Adem relations. This, syntomic Steenrod algebra has been discovered independently in an upcoming work of Shachar Carmeli and Tony Feng [CF], where they use the syntomic power operations to prove the last remaining case of the Tate conjectures on the Artin-Tate pairing on Brauer groups.
@@ 1.3.2. Geometric applications in positive characteristics
    If $X$ is a smooth $k$-scheme and $H^{n}(X)=\mathrm{CH}^{n}(X)\otimes\Lambda$ is the Chow group of codimension $n$-cycles with coefficients in a ring $\Lambda$, one asks if a cycle can be represented by a linear combination of cycles whose components are nonsingular; this is an algebro-gemetric counterpart to the topological question of Borel-Haefliger [BH61]. This was resolved in the negative by [HRT74] over the complex numbers and with integral coefficients. The problem has received renewed attention in recent years when $\Lambda=\mathbb{Q}$ due to the seminal work of Kollar and Voisin [KV24] which proves a positive result for cycles below the middle dimension.
@@
    Let $X$ be a smooth projective variety over the complex numbers or a finitely generated field. Over the complex numbers we have the Hodge conjecture, in which case $H^{n}=H^{n,n}$ the group of Hodge cycles. Otherwise, over certain fields, we have the Tate conjecture in which case $H^{n}$ is the group of Tate cycles, defined as the subgroup of geometric $\ell$-adic cohomology which are fixed by the Galois group. These deep conjectures ask that these groups are spanned by algebraic cycles. There is an industry of counterexamples to the integral versions of these conjectures, beginning with Atiyah-Hirzebruch’s counterexamples [AH62]; we also note Schoen’s formulation of the integral version of the Tate conjectures in [Sch98] and the $\ell$-adic analog of Atiyah–Hirzebruch’s counterexamples by Colliot-Thélène–Szamuely [CTS10].
In §4 we investigate mod-$p$ variants of all of the above questions. We provide integral counterexamples to optimistic versions of these problems based on the construction of the Steenrod operations. We remark that our counterexamples are somewhat different from the $\ell$-adic ones (namely, the ones constructed in [HRT74, Tho54, AH62, CTS10]), though they do build on the geometric insights of these works. In particular our counterexamples highlight some of the new phenomena that occurs when working with $p$-adic coefficients.
@@ Acknowledgments
Lastly, we would also like to thank Tony Feng for discussions about his work on Steenrod operation and Tate’s conjectures on the Brauer group, Jacob Lurie for bringing to our attention the relevance of infinite/deep ramification, Matthew Morrow for his inspiring vision on motivic cohomology which led to the techniques of this paper, Alexander Petrov for discussions around Steenrod operations in algebraic geometry, the Tate conjectures and help with parts of Theorem 4.29, and Burt Totaro for discussions around the integral Tate conjecture.
@@ 2.1.1. The purity transformation
[…]
@@ 2.2. Motivic cohomology
Let $R$ be a coefficient ring, we denote by $HR^{\mathbb{A}^{1}}:=H\mathbb{Z}^{\mathbb{A}^{1}}\otimes R$ where $R$ is regarded as an Eilenberg-Maclane spectrum and the tensor product is given by the action of spectra on $\mathrm{SH}_{\mathbb{Z}}$ in presentable, stable $\infty$-categories. Concretely, we have equivalences, natural in $X$:
@@
[…]
@@
We now explain how the work of Levine in [Lev01], namely Theorem 2.11, essentially establishes this result in cases that we need for this paper. The only point to address is to compare the pushforward map involved in (9) versus the purity map. The key calculation is performed in Appendix C. We also remark that [BEM, Theorem 4.43] provides a different perspective on the localization sequence of Levine via a prismatic approach.
@@
To proceed, we make some reductions. Let $B=\mathrm{Spec}(A)$ and $Z=\mathrm{Spec}(A/\pi)$ and let $\iota:Z\hookrightarrow B$ be the closed immersion and $j:X\setminus Z\hookrightarrow X$ be the complementary open immersion. By tor-independent base change for the purity transformation [DJK18, Proposition 3.2.8], it suffices to prove the result for $X=B$ itself. Therefore we are reduced to checking that $\mathfrak{p}_{\iota}$ is an equivalence. Now, Levine’s Theorem 2.11 implies [^4] Strictly speaking, Theorem 2.11 is a statement about graded sheaves of spectra, not about motivic spectra. However, as the sequence of Eq. (9) is linear over the motivic cohomology of $X$, it can be enhanced into a cofibre sequence in $\mathrm{SH}_{X}$ using e.g. the orientation of the motivic cohomology of $X$. See [AHI24b, Construction 6.5] for an argument outlining how graded sheaves with orientations can be enhanced into motivic spectra. the existence of an equivalence $\iota^{!}H\mathbb{Z}_{B}^{\mathbb{A}^{1}}\simeq H\mathbb{Z}^{\mathbb{A}^{1}}_{Z}\otimes\mathbb{T}^{-1}$; for example combine [Spi18, Theorem 7.4 & 7.18] and Remark 2.9. Therefore, the purity transformation is then a self-map: $\mathfrak{p}_{\iota}:H\mathbb{Z}^{\mathbb{A}^{1}}_{Z}\otimes\mathbb{T}^{-1}\rightarrow H\mathbb{Z}^{\mathbb{A}^{1}}_{Z}\otimes\mathbb{T}^{-1}$. But by the calculation in Proposition C.2 in the appendix this map classifies the fundamental class of $Z$, and therefore by linearity $\mathfrak{p}_{\iota}$ is given by the identity on the level of motivic spectra. ∎

### v2

@@ 1. Operations? Again?
At this point, the algebraic geometer might be rightly annoyed — it seems that all the topologist does is to spoil algebro-geometric dreams using these hieroglyphic-like symbols. Then comes Voevodsky. The motivic Steenrod operations on mod-$\ell$ motivic cohomology in characteristic zero [Voe03b] lie at the heart of the proof of the Bloch-Kato conjectures [Voe03a, Voe11]. In fact, the shape of these operations is at the heart of many issues of a good theory of motives and motivic homotopy. Another notable result is Feng’s thesis [Fen20] where he used Steenrod operations on étale cohomology to settle a conjecture of Tate’s on the Artin-Tate pairing on Brauer groups away from the residue characteristic.
1.1. Operations and the statement of the main theorem
Let us now explain the format of motivic cohomology operations. Let $p$ be a prime, $F$ a field. The formation of mod-$p$ motivic cohomology, assembles into a functor valued in bigraded commutative rings (with the cup product $\cup$ as the multiplicative structure):
$H^{\star}_{\mathrm{mot}}(-,\mathbb{F}_{p}(\star))\colon\mathrm{Sm}_{F}^{\mathrm{op}}\to\mathrm{BiGrCAlg}.$
There is a canonical class
$c\in H_{\mathrm{mot}}^{2}(\mathbb{P}_{F}^{1},\mathbb{F}_{p}(1))(\cong\mathrm{Pic}(\mathbb{P}_{F}^{1})/p),$
classifying the mod-$p$ reduction of the first chern class of $\mathcal{O}(1)$. A bistable cohomology operation (of degree $(r,s)$) [Voe03b, (2.13) & Proposition 2.16] is a collection of natural transformation
$P_{ij}:H^{i}_{\mathrm{mot}}(-,\mathbb{F}_{p}(j))\rightarrow H^{i+r}_{\mathrm{mot}}(-,\mathbb{F}_{p}(j+s)),$
which commutes with multiplication by $c$:
$P_{ij}(x\cup c)=P_{i+2,j+1}(x)\cup c.$
This definition is a direct analog of the definition of stable cohomology operations for singular cohomology. The main goal of this paper is to construct certain canonical, non-trivial bistable operations when $F$ is a field of characteristic $p>0$. These operations are even constructed on the extension of mod-$p$ motivic cohomology to all noetherian $\mathbb{F}_{p}$-schemes, defined and studied by the second author and Morrow [EM23], as well as Kelly and Saito [KS24]. The following is our main theorem.
Theorem 1.1.
Let $p$ be a prime and $F$ be a prime field (in other words, either $\mathbb{Q}$ or $\mathbb{F}_{p}$). There exists natural transformations of presheaves of motivic cohomology groups on Noetherian $\mathbb{F}_{p}$-schemes:
$\mathrm{P}^{n}_{ij}:H^{i}_{\mathrm{mot}}(-,\mathbb{F}_{p}(j))\to H^{i+2n(p-1)}_{\mathrm{mot}}(-,\mathbb{F}_{p}(j+n(p-1)))$
and
$B_{ij}^{n}\colon H_{\mathrm{mot}}^{i}(-;\mathbb{F}_{p}(j))\to H_{\mathrm{mot}}^{i+2n(p-1)+1}(-;\mathbb{F}_{p}(j+n(p-1)).$
Moreover, these operations satisfy the expected properties, namely:
  1. (1)
    $\mathrm{P}^{0}=\mathrm{Id}$ and $\mathrm{B}^{i}=\beta\mathrm{P}^{i}$;
  2. (2)
    the motivic Adem relations hold;
  3. (3)
    the motivic Cartan formulas hold;
  4. (4)
    if $x\in H_{\mathrm{mot}}^{2i}(X;\mathbb{F}_{p}(i))$, then $\mathrm{P}^{i}(x)=x^{p}$;
  5. (5)
    if $y\in H^{j}_{\mathrm{mot}}(X;\mathbb{F}_{p}(k))$ is such that jk<ij-k<i and $k\leq i$, then $\mathrm{P}^{i}(y)=0$;
  6. (6)
We refer the reader to Theorem 3.7 for the motivic Adem and Cartain relations.
@@ 1.2. Operations and endomorphisms
We now explain our theorem in the context of motivic homotopy theory. Let $H\mathbb{F}_{p,F}$ be the motivic spectrum representing mod-$p$ motivic cohomology in Morel–Voevodsky’s category $\mathrm{SH}_{F}$ of motivic spectra over $F$; this motivic spectrum could be built as Bloch’s cycle complex [Blo86] or via Voevodsky’s theory of finite correspondences [VSF00]. As explained in the introduction to [HKO17], there are three related, but a priori different algebra of bigraded operations in motivic homotopy theory summarized in the following zig-zag:
@@
    $\mathcal{M}_{F}^{\star,\star}$ is the algebra of bistable operations in motivic cohomology as explained above in §1.1.
@@
    Finally, $\mathcal{A}_{F}^{\star,\star}$ is, prior to this paper, far only defined if $p$ is invertible in $F$. It is the subalgebra of $\mathcal{M}_{F}^{\star,\star}$ generated by power operations as constructed by Voevodsky in [Voe03b] [^1] Of note are also Brosnan’s Steenrod operations for Chow groups [Bro03] built in a similar manner to Voevodsky. ; nowadays we know that these operations arise from the norm structure on motivic cohomology in the sense of [BH21a]; see [BH21a, Example 7.25] for details.
@@
Ultimately, the construction of power operations rely on the classifying stack $B\Sigma_{n}$ where $\Sigma_{n}$ is the symmetric group on $n$-letters. In effect, Voevodsky in [Voe03b] constructs the $n$-th total power operation as the dotted arrow in the following diagram of spectra [^2] We employ the following notation: for any scheme $S$ we have an adjunction $\sigma^{\infty}:\mathrm{Sh}_{\mathrm{Nis},\mathbb{A}^{1}}(\mathrm{Sm}_{S};\mathrm{Sp})\rightleftarrows\mathrm{SH}_{S}:\omega^{\infty}$. The domain of $\sigma^{\infty}$ is the $\infty$-category of $\mathbb{A}^{1}$-invariant Nisnevich sheaves, also commonly referred to as $S^{1}$-spectra” in the literature. We will also use the further adjunction $\Sigma^{\infty}_{\mathbb{T}}:\mathrm{Sh}_{\mathrm{Nis},\mathbb{A}^{1}}(\mathrm{Sm}_{S};\mathrm{Ani})_{\star}\rightleftarrows\mathrm{SH}_{S}:\Omega^{\infty}_{\mathbb{T}},$ between pointed $\mathbb{A}^{1}$-invariant Nisnevich sheaves of anima and $\mathrm{SH}_{S}$.
@@
The problem is that $BC_{p}$ is $\mathbb{A}^{1}$-contractible in characteristic $p>0$, hence one cannot proceed in this manner [BEH+21, Proposition 5.5]. [^3] Unless otherwise specified, all quotient stacks are taken in the fppf-topology. In principle, one needs to replace $BC_{p}$ by the stack $B\mu_{p}$ and one wants a “total $\mu_{p}$-power operation”: $\omega^{\infty}H\mathbb{F}_{p}(X)\rightarrow\omega^{\infty}H\mathbb{F}_{p}(X\times B\mu_{p})$ out of which one can extract elements of $\mathcal{A}^{\star,\star}$. One approach is to endow a “flat norm” structure on $H\mathbb{F}_{p}$, extending the étale norms of [BH21a]. We do not yet have access to this technology.
@@
In particular, $H_{\mathrm{mot}}^{0}(\mathbb{F}_{p},\mathbb{F}_{p}(\star))=0$ unless $\star=0$. Part of the content of Theorem 1.2 is that even though $\widetilde{\tau}_{p}$-multiples of the Bockstein and power operations are killed by (1), those without such multiples survive and retains their independence relations.
@@
1.3. Relations to [Pri20] and [FS25]
The goal of this paper is to construct power operations on mod-$p$ motivic cohomology of $\mathbb{F}_{p}$-schemes, satisfying the analogs of the relations in topology and mod-$\ell$ motivic cohomology. Our approach is to directly construct a graded-$\mathbb{E}_{1}$-map
@@ 1.3. Relations to [Pri20] and [FS25]
The work of Primozic [Pri20], relying on the work of Frankland–Spitzweck [FS25] constructs power operations $\mathrm{P}^{i}:H_{\mathrm{mot}}^{m}(-;\mathbb{F}_{p}(n))\rightarrow H_{\mathrm{mot}}^{m+2i(p-1)}(-;\mathbb{F}_{p}(n+i(p-1))$ which satisfy the Adem and Cartan formulas on the Chow line diagonal. Let us explain how this is done.
@@
Over fields where $p$ is invertible, this map is an equivalence by [Voe03b, HKO17], but this remains unknown over characteristic $p>0$ fields. Since we do not know how exactly the dual Steenrod algebra looks like, we cannot really define the power operations by (re)taking duals. Nonetheless, the main result of [FS25] shows that $\Psi$ is a split monomorphism over any field. This suggests that there is some hope in producing power operations by taking duals.
@@ 1.4.1. Syntomic realizations
where $H\mathbb{F}_{p}^{\mathrm{syn}}$ denotes the motivic spectrum that represents the mod-$p$ syntomic cohomology (see Proposition 4.14). From this, we obtain mod-$p$ power operations $\mathrm{P}^{i}$ acting on the mod-$p$ syntomic cohomology, and in particular an algebra $\mathcal{A}^{\star,\star}_{\mathrm{syn}}$ acting on $H\mathbb{F}_{p}^{\mathrm{syn}}$ that is generated by the $\mathrm{P}^{i}$ and the Bockstein $\beta$ modulo the Adem relations. This, syntomic Steenrod algebra has been discovered independently in an upcoming work of Shachar Carmeli and Tony Feng [CF25], where they use the syntomic power operations to prove the last remaining case of the Tate conjectures on the Artin-Tate pairing on Brauer groups.
@@ 1.4.2. Geometric applications in positive characteristics
    If $X$ is a smooth $k$-scheme and $H^{n}(X)=\mathrm{CH}^{n}(X)\otimes\Lambda$ is the Chow group of codimension $n$-cycles with coefficients in a ring $\Lambda$, one asks if a cycle can be represented by a linear combination of cycles whose components are nonsingular; this is an algebro-gemetric counterpart to the topological question of Borel–Haefliger [BH61]. This was resolved in the negative by [HRT74] over the complex numbers and with integral coefficients. The problem has received renewed attention in recent years when $\Lambda=\mathbb{Q}$ due to the seminal work of Kollar and Voisin [KV24] which proves a positive result for cycles below the middle dimension.
@@
    Let $X$ be a smooth projective variety over the complex numbers or a finitely generated field. Over the complex numbers we have the Hodge conjecture, in which case $H^{n}=H^{n,n}$ the group of Hodge cycles. Otherwise, over certain fields, we have the Tate conjecture in which case $H^{n}$ is the group of Tate cycles, defined as the subgroup of geometric $\ell$-adic cohomology which are fixed by the Galois group. These deep conjectures ask that these groups are spanned by algebraic cycles. There is an industry of counterexamples to the integral versions of these conjectures, beginning with Atiyah–Hirzebruch’s counterexamples [AH62]; we also note Schoen’s formulation of the integral version of the Tate conjectures in [Sch98] and the $\ell$-adic analog of Atiyah–Hirzebruch’s counterexamples by Colliot-Thélène–Szamuely [CTS10].
[…]
@@ Acknowledgments
Lastly, we would also like to thank Tony Feng for discussions about his work on Steenrod operation and Tate’s conjectures on the Brauer group, Jacob Lurie for bringing to our attention the relevance of infinite/deep ramification, Matthew Morrow for his inspiring vision on motivic cohomology which led to the techniques of this paper, Alexander Petrov for discussions around Steenrod operations in algebraic geometry, the Tate conjectures and help with parts of Theorem 4.34, and Burt Totaro for discussions around the integral Tate conjecture. We also thanks Ben Williams and Marco D’Addezio for useful comments on the previous version of the draft.
@@ 2.1.1. The purity transformation
Recall that if $f:X\rightarrow Y$ is a morphism of schemes, then we say that $f$ is smoothable if it factors as $X\xrightarrow{i}V\xrightarrow{\pi}Y$ where $i$ is a closed immersion and $\pi$ is a smooth morphism. The work of Déglise, Jin and Khan [DJK18] constructs the purity transformaton for any smoothable, lci morphism $f:X\rightarrow Y$
@@ 2.2. Motivic cohomology
Let $R$ be a coefficient ring, we denote by $HR^{\mathbb{A}^{1}}:=H\mathbb{Z}^{\mathbb{A}^{1}}\otimes R$ where $R$ is regarded as an Eilenberg–Maclane spectrum and the tensor product is given by the action of spectra on $\mathrm{SH}_{\mathbb{Z}}$ in presentable, stable $\infty$-categories. Concretely, we have equivalences, natural in $X$:
@@
$(s^{0}1)_{X}$$s^{0}(1_{X})$$(s^{0}\mathrm{KGL})_{X}$$s^{0}(\mathrm{KGL}_{X}).$$\scriptstyle{\simeq}\scriptstyle{\simeq}\scriptstyle{\simeq}$$\scriptstyle{\simeq}$
@@
We now explain how the work of Levine in [Lev01], namely Theorem 2.11, essentially establishes this result in cases that we need for this paper. The only point to address is to compare the pushforward map involved in (9) versus the purity map. The key calculation is performed in Appendix C, but also follows from [FS25, Corollary 8.4]. We also remark that [BEM, Theorem 4.43] provides a different perspective on the localization sequence of Levine via a prismatic approach.
@@
To proceed, we make some reductions. Let $B=\mathrm{Spec}(A)$ and $Z=\mathrm{Spec}(A/\pi)$ and let $\iota:Z\hookrightarrow B$ be the closed immersion and $j:X\setminus Z\hookrightarrow X$ be the complementary open immersion. By tor-independent base change for the purity transformation [DJK18, Proposition 3.2.8], it suffices to prove the result for $X=B$ itself. Therefore we are reduced to checking that $\mathfrak{p}_{\iota}$ is an equivalence. Now, Levine’s Theorem 2.11 implies [^5] Strictly speaking, Theorem 2.11 is a statement about graded sheaves of spectra, not about motivic spectra. However, as the sequence of Eq. (9) is linear over the motivic cohomology of $X$, it can be enhanced into a cofibre sequence in $\mathrm{SH}_{X}$ using e.g. the orientation of the motivic cohomology of $X$. See [AHI24b, Construction 6.5] for an argument outlining how graded sheaves with orientations can be enhanced into motivic spectra. the existence of an equivalence $\iota^{!}H\mathbb{Z}_{B}^{\mathbb{A}^{1}}\simeq H\mathbb{Z}^{\mathbb{A}^{1}}_{Z}\otimes\mathbb{T}^{\otimes-1}$; for example combine [Spi18, Theorem 7.4 & 7.18] and Remark 2.9. Therefore, the purity transformation is then a self-map: $\mathfrak{p}_{\iota}:H\mathbb{Z}^{\mathbb{A}^{1}}_{Z}\otimes\mathbb{T}^{\otimes-1}\rightarrow H\mathbb{Z}^{\mathbb{A}^{1}}_{Z}\otimes\mathbb{T}^{\otimes-1}$. But by the calculation in Proposition C.2 in the appendix this map classifies the fundamental class of $Z$, and therefore by linearity $\mathfrak{p}_{\iota}$ is given by the identity on the level of motivic spectra. ∎